OpenStudy (anonymous):

find the slope of the line that goes through the points (1,18)(-8,121)

OpenStudy (the_fizicx99):

M = (y2-y1) / (x2-x1)

OpenStudy (the_fizicx99):

Plug in your values

OpenStudy (solomonzelman):

\[\huge\color{green}{\frac{y_1-y_2}{x_1-x_2}}\] \[\huge\color{green}{\frac{(18)-(121)}{(1)-(-8)}}\] Do it.

OpenStudy (the_fizicx99):

121 - 18 / -8 - 1 108 / -9 = Solve for that

OpenStudy (solomonzelman):

@tHe_FiZiCx99, check your work.

OpenStudy (solomonzelman):

@tHe_FiZiCx99, My bad, sorry, I see what you did, you plugged in the second values first.

OpenStudy (the_fizicx99):

@SolomonZelman Nope?

OpenStudy (solomonzelman):

You don;t know what I meant.

OpenStudy (the_fizicx99):

I do.

OpenStudy (solomonzelman):

What?

OpenStudy (the_fizicx99):

Both way work, I just prefer using the second value's first.

OpenStudy (solomonzelman):

yes, and I said that used the second values first, (apologized, thinking you made a mistake when you didn't as I noticed.)

OpenStudy (the_fizicx99):

Np :>